OpenStudy (anonymous):

Solve –2x^2 +3x – 9 = 0 using the quadratic formula or completing the square method.

OpenStudy (mathteacher1729):

Do you have a preferred method?

OpenStudy (saifoo.khan):

Try quadratic formula. it's easy.

OpenStudy (anonymous):

whichever method is good.

OpenStudy (saifoo.khan):

all 3 methods will give you the same answer.

OpenStudy (anonymous):

completing the square method requires leading coefficent to be 1, and it works nicely if the "middle coefficient" is even. if not, use the quadratic formula (carefully) and be happy

OpenStudy (anonymous):

\[x= -b \pm \sqrt{b ^{2}-4ac} /2a\] my teacher taught is a song .. its like the happy birthday song x is equal to negative b plus or minus the square root of b squared minus 4 a c all over 2 a

OpenStudy (anonymous):

Does anyone know the answer??

OpenStudy (anonymous):

http://au.answers.yahoo.com/question/index?qid=20100313194503AASfWpU that gives u the answer

OpenStudy (mathteacher1729):

How far did you get with this problem before you got stuck? Like... did you plug the numbers into the quadratic formula... then could not simplify them? Or did you not know how to substitute the appropriate numbers into the quad. formula? etc.

OpenStudy (anonymous):

Jadadenay thanks for helping, but unfortunately that website doens't give me the correct answers.

OpenStudy (anonymous):

what are the answer choices ?

OpenStudy (anonymous):

Mathteacher1729 I'm not sure how to go about solving this equation at all since the x^2 is negative.

OpenStudy (anonymous):

Jadadenay the answer choices are: a. x=3+-3i√7/4 , b.x=3+-9i√7/4 , c.x=-3+-3i√7/4 , d.x=-3+-9i√7/4

OpenStudy (anonymous):

i believe its c

OpenStudy (mathteacher1729):

@dmartinez130 "since x^2 is negative"... That means that the coefficient of x^2 is -1. In other words, wherever you see "-x^2" you it means "(-1) times x^2". :)